# Bridged Posterior: Optimization, Profile Likelihood and a New Approach of Generalized Bayes

### v1

@@ 2.2 Bridged Posterior Distributions and Posterior Propriety
When $z$ is the unique minimizer, we may remove the first integration from both the numerator and denominator, replacing $\textup{d}z$ by $z$. The above distribution can be viewed as obeying an equality constraint, which acts as a bridge between a probabilistic model and an optimization problem. Therefore, we refer to (3) as a “bridged posterior”. To clarify, the above formulation encompasses the setting where $\lambda=(\lambda_{A},\lambda_{B})$, where only the first part $\lambda_{A}$ influences the minimization of $g$, and $\lambda_{B}$ corresponds to the other parameters.
For such a posterior, it is important to ensure the propriety of $\Pi(\lambda\mid y)$. To be rigorous, we note that the derived $L(y,z;\lambda)$ may be a kernel function proportional to a complete density of $y$ up to some missing normalizing constant. Thus it suffices to choose an appropriate $\pi_{0}(\lambda)$ ensuring that $\int\Pi(\textup{d}\lambda\mid y)<\infty$.
A challenge arises when checking integrability (such as when verifying posterior propriety), when $L(y,z;\lambda)$ is intractable due to the lack of a closed-form solution $z$. Generally speaking, mathematically verifying integrability may vary from case to case; we develop a useful strategy in the case when $L(y,z;\lambda)$ is a profile likelihood. Consider the following form
@@
where $z=\arg\min_{\zeta}g(\zeta,y;\lambda)$. In the optimization literature, $\min_{\zeta}g(\zeta,y;\lambda)$ is often referred to as the primal problem, and $z$ as the primal solution. A useful technique is to instead solve an associated dual problem $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$, where $\alpha\in\mathbb{R}^{q}$ is known as the dual variable. For example, the Fenchel dual for convex $g$ is based on the conjugate function $g^{\dagger}(\alpha,y;\lambda):=\sup_{\zeta}\{\alpha^{\top}\zeta-g(\zeta,y;\lambda)\}$, and the Lagrangian dual for $\alpha$ under constraints $\tilde{c}(\alpha)\leq\vec{0}$, where $\tilde{c}(\alpha)\in\mathbb{R}^{q}$ and the inequality holds pointwise, is $g^{\dagger}(\alpha,y;\lambda):=\inf_{\zeta\in\mathbb{R}^{p}}g(\zeta,y;\lambda)+\alpha^{\top}\tilde{c}(\zeta)$, where the dual variable $\alpha\geq\vec{0}$. In many cases, the dual problem may be easier to solve, and with helpful techniques such as variable splitting (discussed further below in the context of latent normal models), one can often obtain $g^{\dagger}(\alpha,y;\lambda)$ in a closed form, even when $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$ may not have a closed form.
Duality theory provides a simple way to produce a useful bound:
@@
For a likelihood taking the form of (4), consider ${\inf}_{\zeta}g(\zeta,y;\lambda)$ as the primal problem, and $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$ as the dual problem with $E$ the feasible region of $\alpha$. Assume that weak duality holds: that is, $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)\leq{\inf}_{\zeta}g(\zeta,y;\lambda)$ for any $\lambda$ in the domain, given $y$. If there exists $\tilde{\alpha}\in E$ such that $\int\exp[-h(y,\lambda)]\exp[-g^{\dagger}(\tilde{\alpha},y;\lambda)]\pi_{0}(\textup{d}\lambda)<\infty$, then $\int\Pi(\textup{d}\lambda\mid y)<\infty$.
@@
The proof follows rather trivially from the fact that $g^{\dagger}(\tilde{\alpha},y;\lambda)\leq\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$. On the other hand, this result leads to a very useful method for checking integrability—we do not have to solve for the optimal dual variable $\hat{\alpha}$ at which $\sup g^{\dagger}(\alpha,y;\lambda)$ is attained. Instead, we just need to find any $\tilde{\alpha}\in E$ that makes the product integrable. Moreover, the criteria for weak duality are straightforward to check: for Fenchel duals, $g$ needs to be convex, and for Lagrangian duals, $g$ can be convex or non-convex.
We now illustrate the application of the theorem via a working example.
@@
We consider a modification of the canonical latent normal model that takes a full likelihood denoted by $\tilde{L}$:
@@
where $v$ is commonly a log-concave density of $y_{i}$ conditionally independent for $i=1,\ldots,n$, $Q(\lambda;x)$ is parameterized by a covariance kernel such as $Q(\lambda;x)_{i,j}=\tau\exp(-\|x_{i}-x_{j}\|^{2}/2b)$ with $x_{i}\in\mathbb{R}^{\tilde{d}}$ the observed predictor/location, and parameter $\lambda=(\tau,b)\in\mathbb{R}^{2}$. In our example, we focus on binary $y_{i}$ from Bernoulli distribution under logistic link $v(y_{i}\mid w_{i})=\exp(y_{i}\zeta_{i})/\{1+\exp(\zeta_{i})\}$. We now minimize $g(\zeta,y;\lambda)=-\log\tilde{L}(y,\zeta;\lambda)$ over $\zeta\in\mathbb{R}^{n}$ to induce a conditionally deterministic $z$, with profile likelihood:
$L(y,z;\lambda)\propto\exp\biggl{\{}-\frac{1}{2}z^{\top}Q^{-1}(\lambda;x)z\biggr{\}}\bigg{\{}\prod_{i=1}^{n}v(y_{i}\mid z_{i})\bigg{\}},$ (6)
$z={\arg\min}_{\zeta}\{-\log L(y,\zeta;\lambda)\}.$
As $g(\zeta,y;\lambda)$ can be conveniently decomposed into the sum of a quadratic function and a convex function, this lends itself to variable splitting using the constraint $u=\zeta$. With $\alpha\in\mathbb{R}^{n}$ the Lagrange multiplier, we have the Langragian dual
@@
whenever $(\alpha+y)\in(0,1)^{n}$; otherwise the infimum takes $\infty$- ∞. Therefore, we have the dual function:
$g^{\dagger}(\alpha,y;\lambda)=-\frac{1}{2}\alpha^{\top}Q\alpha-\sum_{i=1}^{n}\biggl{\{}(a_{i}+y_{i})\log\frac{\alpha_{i}+y_{i}}{1-(\alpha_{i}+y_{i})}-\log\frac{1}{1-(\alpha_{i}+y_{i})}\biggr{\}},$ (7)
@@
Due to some similarity between the above form and the quadratic exponential model in McCullagh (1994), we refer to (6) as a “latent quadratic exponential model”. It is not hard to see that the above is an integrable upper bound for $y$, since $y_{i}\in\{0,1\}$, and $(\alpha_{i}+y_{i})$ and $1-(\alpha_{i}+y_{i})$ are both bounded above for any given $\alpha$. To find an appropriate prior for $\lambda$, note that the second part does not involve $\lambda$ at any fixed $\alpha$. Hence it suffices to find a prior such that for a feasible $\tilde{\alpha}$:
@@
In particular using $Q(\lambda;x)_{i,j}=\tau\exp(-\|x_{i}-x_{j}\|^{2}/2b)$, the matrix spectral norm $\|Q(\lambda;x)\|_{2}\leq n\tau$. As we may choose any feasible $\tilde{\alpha}$, we take the one with $\tilde{\alpha}_{i}=-(1/n)\mathbb{1}(y_{i}=1)+(1/n)\mathbb{1}(y_{i}=0)$. We can see that $\tilde{\alpha}^{\top}Q(\lambda;x)\tilde{\alpha}\leq\|\tilde{\alpha}\|_{2}^{2}\|Q(\lambda;x)\|_{2}=\tau$. Therefore, to make the posterior proper, it suffices to assign a half-normal prior for $\tau$ proportional to $\exp(-c_{1}\tau^{2})$ with $c_{1}>0$ and with any proper prior on $b>0$.
@@
For the above example, strong duality holds: $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)=\inf_{\zeta}g(\zeta,y;\lambda)$. Therefore, we can use the dual ascent algorithm to find $\hat{\alpha}={\arg\max}_{\alpha:(\alpha+y)\in(0,1)^{n}}g^{\dagger}(\alpha,y;\lambda)$, and then set $z=-Q\hat{\alpha}$. Note that neither the dual function (7) nor its gradient with respect to $\alpha$ requires the inversion $Q^{-1}$, an $O(n^{3})$ operation, so that optimization can be carried out very efficiently. At the same time, $L(y,z;\lambda)$ can be evaluated quickly since $z^{\top}Q^{-1}z=\hat{\alpha}^{\top}Q\hat{\alpha}$. In contrast, the latent normal model would involve matrix inversion and decomposition for sampling latent $\zeta$. We defer the numerical experiments to Section 5.
@@ 2.3 Predictive Distribution
$\Pi\bigl{\{}y_{(n+1):(n+k)}\mid y_{1:n}\bigr{\}}\propto\int L\bigl{\{}y_{(n+1):(n+k)}\mid y_{1:n},\lambda\bigr{\}}\Pi(\textup{d}\lambda\mid y_{1:n})\propto\int\frac{L\{y_{1:(n+k)},\hat{z}(y_{1:(n+k)},\lambda),\lambda\}}{L\{y_{1:n},\hat{z}(y_{1:n},\lambda),\lambda\}}\Pi(\textup{d}\lambda\mid y_{1:n}),$
for which we could take each posterior sample of $\lambda$, and simulate a vector $y_{(n+1):(n+k)}$ with kernel proportional to ${L\{y_{1:(n+k)},\hat{z}(y_{1:(n+k)},\lambda),\lambda\}}$. When we lack a way to directly draw from the joint distribution of $y_{(n+1):(n+k)}$, note that
@@
suggesting that we can simulate $y_{n+j}$ sequentially for $j=1\ldots k$. When $y_{n+j}$ is in a low-dimensional (often one-dimensional) space, we can employ a simple algorithm such as rejection sampling.
Note that all elements in $z=\hat{z}(y_{1:(n+k)},\lambda)\in\mathbb{R}^{p}$ may vary according to $y_{1:(n+k)}$; we emphasize this by using the notation $\hat{z}(y_{1:(n+k)},\lambda)$. This implies that when there is no closed-form solution for $z$, there is an additional burden to compute $\hat{z}(y_{1:(n+j)},\lambda)$ wherever $j$ increments to $j+1$. Fortunately, for the optimization problem
$\hat{z}(y_{1:(n+j+1)},\lambda)={\arg\min}_{\zeta}g(\zeta,y_{1:(n+j+1)};\lambda),$
we can initialize $\zeta$ at the last optimal when predicting $y_{n+j},$ $\hat{z}(y_{1:(n+j)},\lambda)$, and then it usually takes a few iterations of optimization steps to converge to $\hat{z}(y_{1:(n+j+1)},\lambda)$. For advanced problems, there is a large literature on online optimization algorithms (Jadbabaie et al., 2015) that can be employed to efficiently obtain sequential updates.
@@
where $z_{w}\in\mathbb{R}^{\tilde{p}},\zeta_{w}\in\mathbb{R}^{\tilde{p}}$ and $z_{b}\in\mathbb{R},\zeta_{b}\in\mathbb{R}$. We treat $x_{i}$ as fixed, so that the above likelihood is viewed as a discrete distribution for $(y_{1},\ldots,y_{n+k})$. The function $h$ is the hinge loss, which takes value zero when $y_{i}=1$, $\zeta_{w}^{\top}x_{i}+\zeta_{b}\geq 1$, or when $y_{i}=-1$, $\zeta_{w}^{\top}x_{i}+\zeta_{b}\leq-1$. Effectively, the loss function penalizes not only the misclassified points $(x_{i},y_{i}):y_{i}(\zeta_{w}^{\top}x_{i}+\zeta_{b})<0$, but also the points in the band between two boundaries $\{x:-1<\zeta_{w}^{\top}x_{i}+\zeta_{b}<1\}$. The inclusion of $(1/2)\lambda\|z_{w}\|^{2}_{2}$ leads to a maximum distance between the two hyperplanes $\{x:z_{w}^{\top}x+z_{b}=1\}$ and $\{x:z_{w}^{\top}x+z_{b}=-1\}$, under some tolerance to non-zero hinge losses, with tolerance controlled by $\lambda>0$. For comparison, if we were to directly use a Gibbs posterior with likelihood of the form of (8)—that is, without the equality constraint so that $z$ is replaced by $\zeta$, then it would hold that
@@
Now, under our bridged posterior approach, denote the conditional optimum $z=\hat{z}\bigl{(}\{y_{i}\}_{i=1}^{n+k},\lambda;\{x_{i}\}_{i=1}^{n+k}\bigr{)}$. Though a closed-form marginal for (8) is not available, we know from the Lagrangian dual with multiplier $\alpha\in\mathbb{R}^{n+k}$ (Chang and Lin, 2011) that the decision hyperplane $\mathcal{C}_{z}=\{x:z_{w}^{\top}x+z_{b}=0\}$ satisfies

### v2

@@ 2.2 Bridged Posterior Distributions and Posterior Propriety
When $z$ is the unique minimizer, we may remove the first integration from both the numerator and denominator, replacing $\textup{d}z$ by $z$. The above distribution can be viewed as obeying an equality constraint, which acts as a bridge between a probabilistic model and an optimization problem. Therefore, we refer to (4) as a bridged posterior. To clarify, the above formulation encompasses the setting of $\lambda=(\lambda_{A},\lambda_{B})$, where only $\lambda_{A}$ influences the minimization of $g$, and $\lambda_{B}$ corresponds to the other parameters.
[…]
A challenge arises when checking integrability (such as when verifying posterior propriety), when $L(y,z;\lambda)$ is intractable due to the lack of a closed-form solution $z$. Generally speaking, mathematically verification of integrability may vary from case to case; we develop a useful strategy in the case when $L(y,z;\lambda)$ is a profile likelihood. Consider the following
@@
where $z=\arg\min_{\zeta}g(\zeta,y;\lambda)$.
For checking posterior propriety, a common approach begins with finding an envelope function $f$, with $f(\lambda)\geq\pi_{0}(\lambda)L(y,z;\lambda)$ for all $\lambda$ in the parameter space, and then establishes that $f$ is integrable under the choice of $\pi_{0}$. To find such an envelope function, duality is a useful technique. To provide relevant background, we refer to $\min_{\zeta}g(\zeta,y;\lambda)$ as the primal problem, and $z$ as the primal solution. Associated with the primal problem is the dual optimization problem $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$, where $\alpha\in\mathbb{R}^{q}$ is the dual variable. For example, the Fenchel dual for convex $g$ is based on the conjugate function $g^{\dagger}(\alpha,y;\lambda):=\sup_{\zeta}\{\alpha^{\top}\zeta-g(\zeta,y;\lambda)\}$, and the Lagrangian dual for $\alpha$ under constraints $\tilde{c}(\alpha)\leq\vec{0}$, where $\tilde{c}(\alpha)\in\mathbb{R}^{q}$ and the inequality holds pointwise, is $g^{\dagger}(\alpha,y;\lambda):=\inf_{\zeta\in\mathbb{R}^{p}}g(\zeta,y;\lambda)+\alpha^{\top}\tilde{c}(\zeta)$, where the dual variable $\alpha\geq\vec{0}$.
The dual function $g^{\dagger}(\alpha,y;\lambda)$ is particularly useful here because of the weak duality. That is, $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)\leq{\inf}_{\zeta}g(\zeta,y;\lambda)$ holds for any $\lambda$ in the feasible region, hence providing convenience for finding the envelope function. We now state the useful bound:
@@
For a likelihood (5), consider ${\inf}_{\zeta}g(\zeta,y;\lambda)$ as the primal problem, and
$\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)$ as the dual problem with $E$ the feasible region of $\alpha$. If there exists $\tilde{\alpha}\in E$ such that $\int\exp[-h(y,\lambda)]\exp[-g^{\dagger}(\tilde{\alpha},y;\lambda)]\pi_{0}(\textup{d}\lambda)<\infty$, then $\int\Pi(\textup{d}\lambda\mid y)<\infty$.
@@
This result leads to a very useful method for checking integrability—we do not have to solve for the optimal dual variable $\hat{\alpha}$ at which $\sup g^{\dagger}(\alpha,y;\lambda)$ is attained. Instead, we just need to find any $\tilde{\alpha}\in E$ that makes the product integrable. Moreover, the criteria for weak duality are straightforward to check: for Fenchel duals, $g$ needs to be convex, and for Lagrangian duals, $g$ can be convex or non-convex. We now illustrate the application of the theorem via a working example.
@@
We modify the canonical latent normal model that uses a full likelihood:
@@
where $v$ is commonly a log-concave density of $y_{i}$ conditionally independent for $i=1,\ldots,n$, $Q(\lambda;x)$ is parameterized by a covariance kernel such as $Q(\lambda;x)_{i,j}=\tau\exp(-\|x_{i}-x_{j}\|^{2}/2b)$ with $x_{i}\in\mathbb{R}^{\tilde{d}}$ the observed predictor/location, and parameter $\lambda=(\tau,b)\in\mathbb{R}^{2}$. In our example, we focus on binary $y_{i}$ from Bernoulli distribution under logistic link $v(y_{i}\mid\zeta_{i})=\exp(y_{i}\zeta_{i})/\{1+\exp(\zeta_{i})\}$. We now minimize $g(\zeta,y;\lambda)=-\log\tilde{L}(y,\zeta;\lambda)$ over $\zeta\in\mathbb{R}^{n}$ to induce a conditionally deterministic $z$, with profile likelihood:
$L(y,z;\lambda)\propto\exp\biggl{\{}-\frac{1}{2}z^{\top}Q^{-1}(\lambda;x)z\biggr{\}}\bigg{\{}\prod_{i=1}^{n}v(y_{i}\mid z_{i})\bigg{\}},$ $z={\arg\min}_{\zeta}\{-\log L(y,\zeta;\lambda)\}.$ (7)
We first show how to find the dual function. As $g(\zeta,y;\lambda)$ can be conveniently decomposed into the sum of a quadratic function and a convex function, this lends itself to variable splitting using the constraint $u=\zeta$. With $\alpha\in\mathbb{R}^{n}$ the Lagrange multiplier, we obtain the Lagrangian dual
@@
whenever $(\alpha+y)\in(0,1)^{n}$; otherwise the infimum is $\infty$- ∞. We have the dual function:
$g^{\dagger}(\alpha,y;\lambda)=-\frac{1}{2}\alpha^{\top}Q\alpha-\sum_{i=1}^{n}\biggl{\{}(a_{i}+y_{i})\log\frac{\alpha_{i}+y_{i}}{1-(\alpha_{i}+y_{i})}-\log\frac{1}{1-(\alpha_{i}+y_{i})}\biggr{\}},$
@@
Due to some similarity between the above form and the quadratic exponential model in McCullagh (1994), we refer to (LABEL:eq:latent_quadratic) as a latent quadratic exponential model.
We now show how to apply Theorem 1 to verify the posterior propriety. We can see that the above is an integrable upper bound for $y$, since $y_{i}\in\{0,1\}$, and both $(\alpha_{i}+y_{i})$ and $1-(\alpha_{i}+y_{i})$ are bounded above. To find an appropriate prior for $\lambda$, the second part does not involve $\lambda$ at any fixed $\alpha$. It suffices to find a prior such that for a feasible $\tilde{\alpha}$:
@@
Using $Q(\lambda;x)_{i,j}=\tau\exp(-\|x_{i}-x_{j}\|^{2}/2b)$, the matrix spectral norm $\|Q(\lambda;x)\|_{2}\leq n\tau$. As we may choose any feasible $\tilde{\alpha}$, we take $\tilde{\alpha}_{i}=-(1/n)\mathbb{1}(y_{i}=1)+(1/n)\mathbb{1}(y_{i}=0)$. Since $\tilde{\alpha}^{\top}Q(\lambda;x)\tilde{\alpha}\leq\|\tilde{\alpha}\|_{2}^{2}\|Q(\lambda;x)\|_{2}=\tau$, it suffices to assign a half-normal prior for $\tau$ proportional to $\exp(-c_{1}\tau^{2})$ with $c_{1}>0$ and with any proper prior on $b>0$.
@@
For the above example, strong duality holds: $\sup_{\alpha}g^{\dagger}(\alpha,y;\lambda)=\inf_{\zeta}g(\zeta,y;\lambda)$. Therefore, we can use the dual ascent algorithm to find $\hat{\alpha}={\arg\max}_{\alpha:(\alpha+y)\in(0,1)^{n}}g^{\dagger}(\alpha,y;\lambda)$, and then set $z=-Q\hat{\alpha}$. Note that neither the dual function (LABEL:eq:dual_cost_lnm) nor its gradient with respect to $\alpha$ requires the inversion $Q^{-1}$, an $O(n^{3})$ operation, so that optimization can be carried out very efficiently. At the same time, $L(y,z;\lambda)$ can be evaluated quickly since $z^{\top}Q^{-1}z=\hat{\alpha}^{\top}Q\hat{\alpha}$. In contrast, the latent normal model would involve matrix inversion and decomposition for sampling latent $\zeta$. We defer the numerical experiments to Section 5.
In general, the function $g$ does not have to be the negative log-likelihood. In the Section E.3, we provide a high-dimensional example of maximal flow problem. The optimization subproblem corresponds to a mechanistic process where flows automatically fill the network to the capacity, whereas the likelihood characterizes the difference between the conditionally optimal flows and observed values.
@@ 2.3 Predictive Distribution
$\Pi\bigl{\{}y_{(n+1):(n+k)}\mid y_{1:n}\bigr{\}}$ $\propto\int L\bigl{\{}y_{(n+1):(n+k)}\mid y_{1:n},\lambda\bigr{\}}\Pi(\textup{d}\lambda\mid y_{1:n})$
$\propto\int\frac{L\{y_{1:(n+k)},\hat{z}(y_{1:(n+k)},\lambda),\lambda\}}{L\{y_{1:n},\hat{z}(y_{1:n},\lambda),\lambda\}}\Pi(\textup{d}\lambda\mid y_{1:n}),$
for which we could take each posterior sample of $\lambda$, and simulate a vector $y_{(n+1):(n+k)}$ with kernel proportional to ${L\{y_{1:(n+k)},\hat{z}(y_{1:(n+k)},\lambda),\lambda\}}$. We use $\propto$ to mean that $\Pi\bigl{\{}y_{(n+1):(n+k)}\mid y_{1:n}\bigr{\}}$ contains the additional normalizing term, so it does integrate to one over $y_{(n+1):(n+k)}$.
When we lack a way to directly draw from the joint distribution of $y_{(n+1):(n+k)}$, note that
@@
suggesting that we can simulate $y_{n+j}$ sequentially for $j=1\ldots k$. When $y_{n+j}$ lies in a low-dimensional (often one-dimensional) space, we can employ a simple algorithm such as rejection sampling. Note that all elements in $z=\hat{z}(y_{1:(n+k)},\lambda)\in\mathbb{R}^{p}$ may vary according to $y_{1:(n+k)}$; we emphasize this by using the notation $\hat{z}(y_{1:(n+k)},\lambda)$. This implies that when there is no closed-form solution for $z$, there is an additional burden to compute $\hat{z}(y_{1:(n+j)},\lambda)$ wherever $j$ increments to $j+1$. Fortunately, for the problem $\hat{z}(y_{1:(n+j+1)},\lambda)={\arg\min}_{\zeta}g(\zeta,y_{1:(n+j+1)};\lambda),$ we can initialize $\zeta$ at the last optimal when predicting $y_{n+j},$ $\hat{z}(y_{1:(n+j)},\lambda)$, and it takes a few iterations of optimization steps to converge to $\hat{z}(y_{1:(n+j+1)},\lambda)$. For advanced problems, there is a large literature on online optimization algorithms (Jadbabaie et al., 2015) that can be employed to efficiently obtain sequential updates.
@@
where $z_{w}\in\mathbb{R}^{\tilde{p}},\zeta_{w}\in\mathbb{R}^{\tilde{p}}$ and $z_{b}\in\mathbb{R},\zeta_{b}\in\mathbb{R}$. We treat $x_{i}$ as fixed, so that the above likelihood is viewed as a discrete distribution for $(y_{1},\ldots,y_{n+k})$. The function $h$ is the hinge loss, which takes value zero when $y_{i}=1$, $\zeta_{w}^{\top}x_{i}+\zeta_{b}\geq 1$, or when $y_{i}=-1$, $\zeta_{w}^{\top}x_{i}+\zeta_{b}\leq-1$. Effectively, the loss function penalizes not only the misclassified points $(x_{i},y_{i}):y_{i}(\zeta_{w}^{\top}x_{i}+\zeta_{b})<0$, but also the points in the band between two boundaries $\{x:-1<\zeta_{w}^{\top}x_{i}+\zeta_{b}<1\}$. The inclusion of $(1/2)\lambda\|z_{w}\|^{2}_{2}$ leads to a maximum distance between the two hyperplanes $\{x:z_{w}^{\top}x+z_{b}=1\}$ and $\{x:z_{w}^{\top}x+z_{b}=-1\}$, under some tolerance to non-zero hinge losses, with tolerance controlled by $\lambda>0$.
For comparison, if we were to directly use a Gibbs posterior with likelihood of the form of (LABEL:eq:svm)—that is, without the equality constraint so that $z$ is replaced by $\zeta$, then it would hold that
@@
Now, under our bridged posterior approach, denote the conditional optimum $z=\hat{z}\bigl{(}\{y_{i}\}_{i=1}^{n+k},\lambda;\{x_{i}\}_{i=1}^{n+k}\bigr{)}$. Though a closed-form marginal for (LABEL:eq:svm) is not available, we know from the Lagrangian dual with multiplier $\alpha\in\mathbb{R}^{n+k}$ (Chang and Lin, 2011) that the decision hyperplane $\mathcal{C}_{z}=\{x:z_{w}^{\top}x+z_{b}=0\}$ satisfies